use properties of rational exponents to simplify the expression
x^2/5 x^4/5 over x^3/5

Answer by fractalier(6550)

Okay, using the laws of exponents, we add them when we multiply like bases and subtract them when we divide like bases, so we get
(x^(2/5) * x^(4/5)) / x^(3/5) =
x^(6/5) / x^(3/5) =
x^(3/5)